|(4x-2)/3|=2
So I know you then write 4x-2/3=2 OR 4x-2/3=-2. I don't know how to solve this. I was thinking to multiply by 3 but it didn't work. Please help

Answer by lwsshak3(11628)

|(4x-2)/3|=2
two solutions:
(4x-2)/3=2
multiply by 3
4x-2=6
4x=8
x=2
..
-(4x-2)/3=2
(4x-2)/3=-2
multiply by 3
4x-2=-6
4x=-4
x=-1
solutions:
x=2
or
x=-1
..
comment: you set up the problem correctly and multiplying by 3 would have gotten you the correct solutions. I believe if you put (4x-2) in parenthesis, you might have seen to do it correctly.